OpenStudy (anonymous):

The Acculturation Rating Scale for Mexican Americans (ARSMA) is a psychological test developed to measure the degree of Mexican/Spanish versus Anglo/English acculturation of Mexican Americans. The distribution of ARSMA scores in a population used to develop the test was approximately normal, with mean 3.0 and standard deviation 0.8. A further study gave ARSMA to 50 first-generation Mexican Americans. The mean of their scores is x bar = 2.36. If the standard deviation for the first-generation population is also 0.8, give a 95% confidence interval for the mean ARSMA score for first-generation

OpenStudy (anonymous):

Mexican Americans. The mean of their scores is x bar = 2.36. If the standard deviation for the first-generation population is also 0.8, give a 95% confidence interval for the mean ARSMA score for first-generation Mexican Americans.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok so we know xbar = 2.36, sigma = 0.8, mu = 3 (not used here but we might as well list it) and n = 50

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

s = sigma/sqrt(n) s = ???

OpenStudy (anonymous):

s= .8/sqrt50

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

0.11314?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's s s = 0.11314

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

now compute the upper and lower bounds L = xbar - C*s U = xbar + C*s

jimthompson5910 (jim_thompson5910):

in this case, C = 1.96 (if you don't remember) because we're working with a 95% CI

jimthompson5910 (jim_thompson5910):

so xbar = 2.36 C = 1.96 s = 0.11314

OpenStudy (anonymous):

whats c

jimthompson5910 (jim_thompson5910):

C = 1.96

OpenStudy (anonymous):

L =2.36 - 1.96 U = 2.36 plus 1.96

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

good, so L = ?? U = ??

OpenStudy (anonymous):

l=.4

OpenStudy (anonymous):

4.32

OpenStudy (anonymous):

LL: 2.36 minus 1.96 = 0.4 UL: 2.36 plus 1.96 = 4.32

jimthompson5910 (jim_thompson5910):

oh wait...not sure why i didn't catch that typo

jimthompson5910 (jim_thompson5910):

my bad

jimthompson5910 (jim_thompson5910):

L = xbar - C*s U = xbar + C*s

jimthompson5910 (jim_thompson5910):

L = xbar - C*s L = 2.36 - 1.96*0.11314 L = ???

OpenStudy (anonymous):

So its LL= 2.1382456

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

U = xbar + C*s U = 2.36 + 1.96*0.11314 U = ???

OpenStudy (anonymous):

2.5817544

OpenStudy (anonymous):

--> (2.1382456, 2.5817544) is the answer?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

that's all there is to it with confidence intervals

OpenStudy (anonymous):

thanks so much! i have two more proibs, ill repost